# A new parametrization of Hubble function and Hubble tension

### v2

@@ Abstract
We present a new Hubble parameterization method and employ observational data from Hubble, Pantheon, and Baryon Acoustic Oscillations to constrain model parameters. The proposed method is thoroughly validated against these datasets, demonstrating a robust fit to the observational data. The obtained best-fit values are $H_{0}=67.5^{+1.3}_{-1.6}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.2764\pm{0.0094}$, and $\alpha=0.33\pm{0.22}$, consistent with the Planck 2018 results, highlighting the existence of Hubble tension.
@@ I Introduction
Hubble constant ($H_{0}$) is a fundamental parameter that can quantify the rate of cosmic expansion. It is frequently employed to elucidate the motion of celestial bodies, such as galaxies, relative to the observer’s position. In recent decades, measurements of the Hubble constant have garnered attention within the scientific community due to notable discrepancies among results obtained from diverse measurement methodologies and data sources [1]. Early measurements of the Hubble constant often entailed observations of primordial cosmic signals, such as the cosmic microwave background radiation (CMB). For instance, using the standard cosmological model, $\Lambda$CDM, a Hubble constant of $H_{0}=67.4\pm 0.5$ $\text{km s}^{-1}\text{Mpc}^{-1}$ was derived [2]. Integrated Baryon Acoustic Oscillations (BAO) measurements with cosmic microwave background (CMB) data from WMAP resulted in $H_{0}=67.63\pm 1.30$ $\text{km s}^{-1}\text{Mpc}^{-1}$ [3]. Early Hubble constant measurements consistently indicate values lower than 70 $\text{km s}^{-1}\text{Mpc}^{-1}$ [4, 5]. The age of an old quasar APM 08279+5255 at $z=3.91$ also tends to support a lower Hubble constant [6]
@@
Later Hubble constant measurements primarily rely on astrophysical observations such as supernovae and galaxies. A result of $H_{0}=73.04\pm 1.04$ $\text{km s}^{-1}\text{Mpc}^{-1}$ was derived based on the Cepheid-SN (Cepheid-supernova) sample [7]. According to [8], they employed a joint analysis of six strongly lensed gravitational lensing events with measured time delays, providing a Hubble constant estimate of $H_{0}=73.3^{+1.7}_{-1.8}$ $\text{km s}^{-1}\text{Mpc}^{-1}$. The detection of gravitational wave events from neutron star mergers by LIGO and Virgo yielded an estimate of $H_{0}=70^{+12}_{-8}$ $\text{km s}^{-1}\text{Mpc}^{-1}$ in 2017 [9]. Using the Hubble Space Telescope, the Hubble constant was directly measured as $H_{0}=74.03\pm 1.42$ $\text{km s}^{-1}\text{Mpc}^{-1}$ through the distance ladder method, providing calibration for the magnitude-redshift relation for 253 Type Ia supernovae [10]. It is evident that numerous late-time measurements favor $H_{0}>70$ $\text{km s}^{-1}\text{Mpc}^{-1}$ with a minority reporting a value of $H_{0}\approx 70$ $\text{km s}^{-1}\text{Mpc}^{-1}$.
@@
Observations of the early universe, including CMB data, typically yield lower values for the Hubble constant ($H_{0}$). Conversely, measurements of celestial bodies at closer distances, such as supernovae and other large-scale cosmic structures, result in higher values for $H_{0}$. The inconsistency between the results obtained from these two methods has captured the attention of researchers. This disparity is referred to as the Hubble tension [4, 11, 12, 13, 14], and the statistical significance of these differences surpasses the range of measurement errors, prompting significant discussion and research. The existence of Hubble tension suggests the possibility of unknown physical processes or issues in observational systems concerning the evolution and nature of the universe [4, 15, 16, 17, 18]. Efforts to address this issue include improvements in data analysis, reduction of systematic errors, the adoption of new observational methods, and a re-examination of cosmological models.
@@
In addressing the Hubble tension, Lin et al. [19] proposed a potential resolution in the Early Dark Sector (EDS), where dark matter mass depends on the Early Dark Energy (EDE) scalar field. They explored a Plank-suppressed EDE coupled with dark matter, finding that this Triggered Early Dark Sector (tEDS) model naturally resolves the coincidence problem of EDE on the background level. Fitting the current cosmological data, including local distance gradients and low-redshift amplitudes of fluctuations, they obtained a Hubble constant of $H_{0}=71.2$ $\text{km s}^{-1}\text{Mpc}^{-1}$. In the presence of non-standard cosmology, a reconciliation between CMB and local measurements has yielded $H_{0}=70-74$ $\text{km s}^{-1}\text{Mpc}^{-1}$ [20]. According to [21], they explored a novel dark fluid model known as the Exponential Acoustic Dark Energy (eADE) model to alleviate the tension in the Hubble telescope. Comparisons with the standard model resulted in $H_{0}=70.06^{+1.13}_{-1.09}$ $\text{km s}^{-1}\text{Mpc}^{-1}$. Some dynamical dark energy models, see for example [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], could also reduce the Hubble tension.
@@
II A new parametrization of Hubble parameter
According to the Planck 2018 results, the spacetime is spatially flat: $\Omega_{\rm K0}=0.001\pm 0.002$ [2], so here we consider a flat Friedmann-Robertson-Walker-Lemaître (FRWL) spacetime
@@ II A new parametrization of Hubble parameter
where $a(t)$ is the scale factor. We use the unit $c=1$ here. The Friedmann equations take the form
$H^{2}\equiv\left(\frac{\dot{a}}{a}\right)^{2}=H^{2}_{0}\left[\Omega_{\rm{m0}}(1+z)^{3}+\Omega_{\rm{r0}}(1+z)^{4}+\Omega_{\rm{x}}(z)\right],$ (2)
@@
There are usually two way to parameterize dark energy: one way is to parameterize the EoS (such as the widely used CPL parameterization model [45, 46]), the other is to directly parameterize the Hubble function. Here we adopt the latter approach. Since the Hubble tension mainly rises from Planck 2018 (based on $\Lambda$CDM model) and Cepheid calibrated supernovae Ia measurements [47], we consider a parameterization slightly different from $\Lambda$CDM. Furthermore, since the degeneracy between parameters could affect the fitting results, we consider the model with as few parameters as possible. The Hubble founction we suggest takes the following form
@@
That is to say, the contributions of radiation and dark energy are combined into one term.
@@ III Observational data and methodology
In the previous section, we discussed a new Hubble parameterization method, and now we aim to validate whether the approximate values of the model parameters can effectively describe the current universe based on observational data. We primarily employ three datasets, namely the Hubble dataset with 62 data points, the Pantheon dataset with 1701 data points, and a set of six Baryon BAO datasets. For numerical analysis and parameter constraints using the mentioned datasets, we utilize the emcee code [48]. Additionally, to understand the outcomes of the MCMC study, we employ 64 walkers and 2000 steps across all datasets. And the final results will be discussed in the form of 2D contour plots with 1 $\sigma$ and 2 $\sigma$ errors.
@@ III.1 Observational Hubble data
Utilizing Hubble observational data to constrain cosmological models is a significant methodology. This approach involves measuring the historical expansion of the universe to derive constraints on cosmological parameters. The Hubble parameter, denoted as $H(z)$, characterizes the rate of cosmic expansion and serves as a fundamental cosmological quantity. Its dependence on redshift ($z$) provides essential insights into the cosmic evolution [49, 50].
@@
In this part, $dz$ is obtained through spectroscopic surveys, making a measurement of $dt$ a means to derive the model-independent value of the Hubble parameter. The $H(z)$ data set we use consists of 34 $H(z)$ measurements obtained by calculating the differential ages of galaxies, which is called cosmic chronometer [51, 52, 53, 54, 55, 52, 56], and 28 $H(z)$ measurements inferred from the BAO peak in the galaxy power spectrum [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69], as collected in [70]. As the data provided by the DA method is independent of cosmological models, it can be employed to explore alternative cosmological models. The range of these data points is 0 $<z<$ 1.965. Furthermore, we adopt an intermediate value of $H_{0}=70$ km/s/Mpc for our analysis [71]. To determine the mean values of the model parameters $H_{0}$, $\Omega_{\rm{m0}}$ and $\alpha$ (using maximum likelihood analysis), we employ the chi-square function as follows [72]:
@@
In Figure 2, we present error bar plots for the aforementioned Hubble data, compared with the $\Lambda$CDM model ($H_{0}=67.4$ km/s/Mpc and $\Omega_{\rm{m0}}=0.315$) [2], indicating that $\alpha$CDM effectively captures the observed Hubble dataset.
@@ III.2 Observational Pantheon data
SNe Ia commonly known as standard candles, serve as powerful distance probes for studying the cosmological dynamics of the universe. Over the past two decades, the sample size of SNe Ia datasets has steadily increased. We employ the largest supernova Ia sample to date, Pantheon+, which amalgamates data from various surveys such as the Sloan Digital Sky Survey (SDSS), the SNe Legacy Survey (SNLS), the Hubble Space Telescope (HST) survey, and others, comprising 1701 confirmed supernovae from 18 different surveys [73, 7]. The Pantheon+ dataset spans a redshift range of z $\in$ (0.0012, 2.2614), with a notable increase in the number of SNe at low redshifts. We need to fit the model parameters by comparing the theoretical distance modulus $\mu_{\rm{th}}$ values with the observed $\mu_{\rm{obs}}$ values. Each distance modulus can be computed using the following formula:
@@
In this expression, $C_{\rm{SN}}$ represents the covariance matrix, as introduced by Suzuki [74]. Additionally, $\Delta\mu{i}=\mu_{\rm{th}}(z_{i},H_{0},\Omega_{\rm{m0}},\alpha)-\mu_{\rm{obs}}$ is defined as the disparity between the observed distance modulus value, derived from cosmic data, and its theoretical counterpart generated from the model using the parameter space $H_{0}$, $\Omega_{\rm{m0}}$ and $\alpha$.
@@
We also use Hubble and Pantheon data to fit $w$CDM, obtaining $H_{0}=67.2\pm 1.8$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.263\pm{0.011}$, and $\omega_{\rm{x}}=-0.845\pm{0.088}$, respectively. The best-fitting values are listed in Table 2 for comparison. We observe that $\alpha$CDM is also consistent with $w$CDM within 1 $\sigma$ error, allowed by Hubble+Pantheon data.
@@ III.3 Observational Baryon Acoustic Oscillations data
BAO arises from acoustic density waves in the early universe’s primordial plasma, causing fluctuations in the observable baryonic matter density in the cosmos. Detecting BAO involves large-scale surveys and redshift measurements to gather information about the large-scale structure of the universe. BAO detectors offer highly precise measurements of large-scale structures, largely unaffected by uncertainties in the nonlinear evolution of the matter density field and other systematic errors. They are considered a standard ruler for measuring the cosmological background evolution [75]. To improve statistical significance, broaden the redshift range, and gain more comprehensive cosmological insights, we utilize the combined data from six different BAO measurements at various redshifts [76, 77, 78]. The information taken from the BAO peaks in the matter power spectrum can be used to determine the Hubble parameter $H(z)$ and the angular diameter distance $d_{A}(z)$ which takes the form
@@
The combination of the angular diameter distance and the Hubble parameter, $D_{V}(z)$, is determined by [79]
@@
The inverse covariance matrix $C_{\rm{BAO}}^{-1}$ is represented in [78]. The six BAO datasets are provided in Table 3. At $z_{\ast}\simeq 1091$, photon decoupling occurred, allowing the CMB to propagate through the universe, eventually becoming the observed cosmic microwave background radiation today. This redshift value is derived through detailed observations and analysis of the CMB [80].
@@
We also use Hubble+Pantheon+BAO data to constrain $w$CDM, obtaining $H_{0}=67.3\pm{1.4}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.2772\pm{0.0098}$, and $\omega_{\rm{x}}=-0.885\pm 0.072$, respectively. For comparison, we list the best-fitting values in Table 4 and observe that $\alpha$CDM is still consistent with $w$CDM within 1 $\sigma$ error, again implying that the impact of radiation is relatively weak, allowed by Hubble+Pantheon+BAO data.

### v3

@@ Abstract
We present a new parameterized Hubble function and employ observational data from Hubble, Pantheon, and Baryon Acoustic Oscillations to constrain model parameters. The proposed method is thoroughly validated against these datasets, demonstrating a robust fit to the observational data. The obtained best-fit values are $H_{0}=67.5^{+1.3}_{-1.6}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.2764\pm{0.0094}$, and $\alpha=0.33\pm{0.22}$, consistent with the Planck 2018 results, highlighting the existence of Hubble tension.
@@ I Introduction
Hubble constant ($H_{0}$) is a fundamental parameter that can quantify the rate of cosmic expansion. It is frequently employed to elucidate the motion of celestial bodies, such as galaxies, relative to the observer’s position. In recent decades, measurements of the Hubble constant have garnered attention within the scientific community due to notable discrepancies among results obtained from diverse measurement methodologies and data sources da Costa et al. (2023). Early measurements of the Hubble constant often entailed observations of primordial cosmic signals, such as the cosmic microwave background radiation (CMB). For instance, using the standard cosmological model, $\Lambda$CDM, a Hubble constant of $H_{0}=67.4\pm 0.5$ $\text{km s}^{-1}\text{Mpc}^{-1}$ was derived Aghanim et al. (2020). Integrated Baryon Acoustic Oscillations (BAO) measurements with cosmic microwave background (CMB) data from WMAP resulted in $H_{0}=67.63\pm 1.30$ $\text{km s}^{-1}\text{Mpc}^{-1}$ Zhang and Huang (2019). Early Hubble constant measurements consistently indicate values lower than 70 $\text{km s}^{-1}\text{Mpc}^{-1}$ Di Valentino et al. (2021a); Anchordoqui et al. (2021). The age of an old quasar APM 08279+5255 at $z=3.91$ also tends to support a lower Hubble constant Yang and Zhang (2010)
@@
Later Hubble constant measurements primarily rely on astrophysical observations such as supernovae and galaxies. A result of $H_{0}=73.04\pm 1.04$ $\text{km s}^{-1}\text{Mpc}^{-1}$ was derived based on the Cepheid-SN (Cepheid-supernova) sample Riess et al. (2022). According to Wong et al. (2019), they employed a joint analysis of six strongly lensed gravitational lensing events with measured time delays, providing a Hubble constant estimate of $H_{0}=73.3^{+1.7}_{-1.8}$ $\text{km s}^{-1}\text{Mpc}^{-1}$. The detection of gravitational wave events from neutron star mergers by LIGO and Virgo yielded an estimate of $H_{0}=70^{+12}_{-8}$ $\text{km s}^{-1}\text{Mpc}^{-1}$ in 2017 Abbott et al. (2017). Using the Hubble Space Telescope, the Hubble constant was directly measured as $H_{0}=74.03\pm 1.42$ $\text{km s}^{-1}\text{Mpc}^{-1}$ through the distance ladder method, providing calibration for the magnitude-redshift relation for 253 Type Ia supernovae Riess et al. (2011). It is evident that numerous late-time measurements favor $H_{0}>70$ $\text{km s}^{-1}\text{Mpc}^{-1}$ with a minority reporting a value of $H_{0}\approx 70$ $\text{km s}^{-1}\text{Mpc}^{-1}$.
@@
Observations of the early universe, including CMB data, typically yield lower values for the Hubble constant ($H_{0}$). Conversely, measurements of celestial bodies at closer distances, such as supernovae and other large-scale cosmic structures, result in higher values for $H_{0}$. The inconsistency between the results obtained from these two methods has captured the attention of researchers. This disparity is referred to as the Hubble tension (Di Valentino et al., 2021a, b; Wang et al., 2023; Abdalla et al., 2022; Dainotti et al., 2021), and the statistical significance of these differences surpasses the range of measurement errors, prompting significant discussion and research. The existence of Hubble tension suggests the possibility of unknown physical processes or issues in observational systems concerning the evolution and nature of the universe (Di Valentino et al., 2021a; Verde et al., 2019a; Di Valentino et al., 2019; Freedman, 2017; Riess, 2019). Efforts to address this issue include improvements in data analysis, reduction of systematic errors, the adoption of new observational methods, and a re-examination of cosmological models.
@@
In addressing the Hubble tension, Lin et al. Lin et al. (2023) proposed a potential resolution in the Early Dark Sector (EDS), where dark matter mass depends on the Early Dark Energy (EDE) scalar field. They explored a Plank-suppressed EDE coupled with dark matter, finding that this Triggered Early Dark Sector (tEDS) model naturally resolves the coincidence problem of EDE on the background level. Fitting the current cosmological data, including local distance gradients and low-redshift amplitudes of fluctuations, they obtained a Hubble constant of $H_{0}=71.2$ $\text{km s}^{-1}\text{Mpc}^{-1}$. In the presence of non-standard cosmology, a reconciliation between CMB and local measurements has yielded $H_{0}=70-74$ $\text{km s}^{-1}\text{Mpc}^{-1}$ Alcaniz et al. (2022a). According to Yin (2022), they explored a novel dark fluid model known as the Exponential Acoustic Dark Energy (eADE) model to alleviate the tension in the Hubble telescope. Comparisons with the standard model resulted in $H_{0}=70.06^{+1.13}_{-1.09}$ $\text{km s}^{-1}\text{Mpc}^{-1}$. Some dynamical dark energy models, see for example Poulin et al. (2019); Sakstein and Trodden (2020); Karwal et al. (2022); McDonough et al. (2022); Tutusaus et al. (2023); Dahmani et al. (2023); Alonso-López et al. (2023); Montani et al. (2023a); Torres-Arzayus et al. (2023); Li and Shafieloo (2019); Pan et al. (2019); Panpanich et al. (2021); De Felice et al. (2020); Alestas et al. (2020, 2022); Castillo-Santos et al. (2023); Dinda (2022); Sharma et al. (2023); Cai et al. (2021); Montani et al. (2023b); Fang et al. (2024); Liu et al. (2024a, b), could also reduce the Hubble tension.
@@
II A new parametrization of Hubble function
According to the Planck 2018 results, the spacetime is spatially flat: $\Omega_{\rm K0}=0.001\pm 0.002$ Aghanim et al. (2020), so here we consider a flat Friedmann-Robertson-Walker-Lemaître (FRWL) spacetime
@@ II A new parametrization of Hubble function
where $a(t)$ is the scale factor. We use the unit $c=1$ here. For a perfect fluid with energy-momentum tensor, $T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu}$, where $\rho$ and $p$ are the energy density and the pressure in the rest frame of the fluid, the 00 component of the Einstein equation, $G_{\mu\nu}=8\pi GT_{\mu\nu}$, yields the Friedmann equation
$H^{2}\equiv\left(\frac{\dot{a}}{a}\right)^{2}=\frac{8\pi G}{3}\rho_{\rm t},$ (2)
where $\rho_{\rm t}$ represents the total energy density contributed from non-relative matter (dark matter and baryonic matter), radiation, and dark energy. Defining dimensionless density parameters $\Omega_{i0}=\rho_{i0}/\rho_{\rm{c}}$, where the subscript ‘i’ denotes a certain energy component, the subscript ‘0’ denotes quantities evaluated today, and the critic density is $\rho_{\rm{c}}=3H^{2}_{0}/(8\pi G)$, then the Friedmann equation takes the form
$H^{2}=H^{2}_{0}\left[\Omega_{\rm{m0}}(1+z)^{3}+\Omega_{\rm{r0}}(1+z)^{4}+\Omega_{\rm{x}}(z)\right],$ (3)
@@
There are usually two way to parameterize dark energy: one way is to parameterize the EoS (such as the widely used CPL parameterized model Chevallier and Polarski (2001); Linder (2003)), the other is to directly parameterize the Hubble function. Here we adopt the latter approach. Since the Hubble tension mainly rises from Planck 2018 (based on $\Lambda$CDM model) and Cepheid calibrated supernovae Ia measurements Verde et al. (2019b), we consider a parameterized slightly different from $\Lambda$CDM. Furthermore, since the degeneracy between parameters could affect the fitting results, we consider the model with as few parameters as possible. The Hubble function we suggest takes the following form
@@
Comparing with other parameterized models, $\alpha$CDM has only one more parameter than $\Lambda$CDM and unifies the contributions of radiation and dark energy into one term. Due to fewer parameters, $\alpha$CDM could effectively reduce the degeneracies between parameters and so can be used to fit low redshift observational data (because of the limited amount of data, it is difficult to use low redshift observational data to effectively constrain models with more parameters).
@@ III Observational data and methodology
In the previous section, we discussed a new Hubble parameterized method, and now we aim to validate whether the approximate values of the model parameters can effectively describe the current universe based on observational data. We primarily employ three datasets, namely the Hubble dataset with 62 data points, the Pantheon dataset with 1701 data points, and a set of six Baryon BAO datasets. For numerical analysis and parameter constraints using the mentioned datasets,
we employ the MCMC method for the global fitting of model parameters. MCMC is a widely-used and effective technique for cosmological parameter estimation. This method allows for thorough exploration of the parameter space, enabling us to find the optimal parameter values while also estimating their uncertainties. The specific steps are as follows: first we define a cosmological model with parameters such as $H_{0}$, $\Omega_{\rm{m}0}$, and $\alpha$; then by comparing the observed values from each dataset with the theoretical predictions, we construct the likelihood function for each dataset; during the MCMC sampling process, we integrate these datasets by maximizing the likelihood function for Hubble parameter measurements, supernova distance moduli distribution, and BAO peak positions. The use of a combined likelihood function allows us to appropriately weight and combine the error distributions of different datasets, yielding a global likelihood function. Here we utilize the emcee package, which is a Python implementation of MCMC that is especially effective for parallel exploration of multi-dimensional parameter spaces Goodman and Weare (2010). Additionally, to understand the outcomes of the MCMC study, we employ 64 walkers and 2000 steps across all datasets. And the final results will be discussed in the form of 2D contour plots with 1 $\sigma$ and 2 $\sigma$ errors.
@@ III.1 Observational Hubble data
Utilizing Hubble observational data to constrain cosmological models is a significant methodology. This approach involves measuring the historical expansion of the universe to derive constraints on cosmological parameters. The Hubble parameter, denoted as $H(z)$, characterizes the rate of cosmic expansion and serves as a fundamental cosmological quantity. Its dependence on redshift ($z$) provides essential insights into the cosmic evolution Salehi and Hatami (2022); Santos et al. (2016).
@@
In this part, $dz$ is obtained through spectroscopic surveys, making a measurement of $dt$ a means to derive the model-independent value of the Hubble parameter. The $H(z)$ data set we use consists of 34 $H(z)$ measurements obtained by calculating the differential ages of galaxies, which is called cosmic chronometer Zhang et al. (2014); Stern et al. (2010); Moresco et al. (2012, 2016); Ratsimbazafy et al. (2017); Stern et al. (2010); Moresco (2015), and 28 $H(z)$ measurements inferred from the BAO peak in the galaxy power spectrum Gaztanaga et al. (2009); Chuang and Wang (2013); Blake et al. (2012); Busca et al. (2013); Oka et al. (2014); Font-Ribera et al. (2014); Anderson et al. (2014); Delubac et al. (2015); Wang et al. (2017); Alam et al. (2017a); Bautista et al. (2017); Zhao et al. (2019); Borghi et al. (2022), as collected in Yang (2024). As the data provided by the DA method is independent of cosmological models, it can be employed to explore alternative cosmological models. The range of these data points is 0 $<z<$ 1.965. Furthermore, we adopt an intermediate value of $H_{0}=70$ km/s/Mpc for our analysis Perivolaropoulos and Skara (2022). To determine the mean values of the model parameters $H_{0}$, $\Omega_{\rm{m0}}$ and $\alpha$ (using maximum likelihood analysis), we employ the chi-square function as follows Giostri et al. (2012):
@@
In Figure 2, we present error bar plots for the aforementioned Hubble data, compared with the $\Lambda$CDM model ($H_{0}=67.4$ km/s/Mpc and $\Omega_{\rm{m0}}=0.315$) Aghanim et al. (2020), indicating that $\alpha$CDM effectively captures the observed Hubble dataset.
@@ III.2 Observational Pantheon data
SNe Ia commonly known as standard candles, serve as powerful distance probes for studying the cosmological dynamics of the universe. Over the past two decades, the sample size of SNe Ia datasets has steadily increased. We employ the largest SNe Ia sample to date, Pantheon+, which amalgamates data from various surveys such as the Sloan Digital Sky Survey (SDSS), the SNe Legacy Survey (SNLS), the Hubble Space Telescope (HST) survey, and others, comprising 1701 confirmed supernovae from 18 different surveys Brout et al. (2022); Riess et al. (2022). The Pantheon+ dataset spans a redshift range of z $\in$ (0.0012, 2.2614), with a notable increase in the number of SNe at low redshifts. We need to fit the model parameters by comparing the theoretical distance modulus $\mu_{\rm{th}}$ values with the observed $\mu_{\rm{obs}}$ values. Each distance modulus can be computed using the following formula:
@@
In this expression, $C_{\rm{SN}}$ represents the covariance matrix, as introduced by Suzuki Suzuki et al. (2012). Additionally, $\Delta\mu{i}=\mu_{\rm{th}}(z_{i},H_{0},\Omega_{\rm{m0}},\alpha)-\mu_{\rm{obs}}$ is defined as the disparity between the observed distance modulus value, derived from cosmic data, and its theoretical counterpart generated from the model using the parameter space $H_{0}$, $\Omega_{\rm{m0}}$ and $\alpha$.
@@
We also use Hubble and Pantheon data to fit $w$CDM, obtaining $H_{0}=67.2\pm 1.8$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.263\pm{0.011}$, and $w_{\rm{x}}=-0.845\pm{0.088}$, respectively. The best-fitting values are listed in Table 2 for comparison. We observe that $\alpha$CDM is also consistent with $w$CDM within 1 $\sigma$ error, allowed by Hubble+Pantheon data.
@@ III.3 Observational Baryon Acoustic Oscillations data
BAO arises from acoustic density waves in the early universe’s primordial plasma, causing fluctuations in the observable baryonic matter density in the cosmos. Detecting BAO involves large-scale surveys and redshift measurements to gather information about the large-scale structure of the universe. BAO detectors offer highly precise measurements of large-scale structures, largely unaffected by uncertainties in the nonlinear evolution of the matter density field and other systematic errors. They are considered a standard ruler for measuring the cosmological background evolution Scolnic et al. (2018). To improve statistical significance, broaden the redshift range, and gain more comprehensive cosmological insights, we utilize the combined data from six different BAO measurements at various redshifts Beutler et al. (2011); Alam et al. (2017b); Myrzakulov et al. (2023). The information taken from the BAO peaks in the matter power spectrum can be used to determine the Hubble parameter $H(z)$ and the angular diameter distance $d_{A}(z)$ which takes the form
@@
The combination of the angular diameter distance and the Hubble parameter, $D_{V}(z)$, is determined by Eisenstein et al. (2005)
@@
The inverse covariance matrix $C_{\rm{BAO}}^{-1}$ is represented in Myrzakulov et al. (2023). The six BAO datasets are provided in Table 3. At $z_{\ast}\simeq 1091$, photon decoupling occurred, allowing the CMB to propagate through the universe, eventually becoming the observed cosmic microwave background radiation today. This redshift value is derived through detailed observations and analysis of the CMB Alcaniz et al. (2022b).
@@
We also use Hubble+Pantheon+BAO data to constrain $w$CDM, obtaining $H_{0}=67.3\pm{1.4}$ $\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{\rm{m0}}=0.2772\pm{0.0098}$, and $w_{\rm{x}}=-0.885\pm 0.072$, respectively. For comparison, we list the best-fitting values in Table 4 and observe that $\alpha$CDM is still consistent with $w$CDM within 1 $\sigma$ error, again implying that the impact of radiation is relatively weak, allowed by Hubble+Pantheon+BAO data.